OpenStudy (angelab97):

Let f(X) be defines for all x by the equation f(x)= 12x+8. Thus, f(2)=32. If f(x) / f(0)=2x, then find the? value of x. I'm so confused with this problem. Someone please help me solve this problem! Please show work and explain each step because I want to understand how to the problem

OpenStudy (angelab97):

what is the question asking for?

terenzreignz (terenzreignz):

Okay, first of all, we have a defined function \[\Large f(\color{red}x)=12\color{red}x + 8 \] right?

OpenStudy (angelab97):

yes

OpenStudy (angelab97):

the answer is supposed to be 2

terenzreignz (terenzreignz):

Let's find out. Your problem also tells you (though only for the sake of information) that \[\Large f(\color{red}2) = 12(\color{red}2)+ 8 = 24+8 = \color{blue}{32}\]

terenzreignz (terenzreignz):

Now, what is the value of \(\large f(\color{red}0) = \color{green}?\)

OpenStudy (angelab97):

that's what confused me. How do i find that out?

terenzreignz (terenzreignz):

Well, when you have \[\Large f(\color{red}x) = <some \ expression \ involving \ \color{red}x>\] Then, say \[\Large f(\color{red}a) \] just asks you to replace all the x's with a's. Just like in this example, we have \[\Large f(\color{red}x) = 12\color{red}x+8\] \[\Large f(\color{red}2) = 12(\color{red}2) + 8=\color{red}{24}+8 = \color{blue}{32}\]

terenzreignz (terenzreignz):

As you can see, to get the value of \(\large f(\color{red}2)\), we simply replaced the \(\large \color{red}x\) in \(\large 12\color{red}x+8\) with a \(\large \color{red}2\) and then solved :)

OpenStudy (angelab97):

so for f(0) it would be 12(0) +8=8

terenzreignz (terenzreignz):

YES. Exactly :)

terenzreignz (terenzreignz):

So, we are given.. \[\Large \frac{f(\color{red}x)}{f(\color{red}0)}=2\color{red}x\]

terenzreignz (terenzreignz):

\(\large f(\color{red}x)\) is just \(\large 12\color{red}x+8\) so we replace it with that... \[\Large \frac{12\color{red}x+8}{f(\color{red}0)}= 2\color{red}x\]

terenzreignz (terenzreignz):

What's \(\large f(\color{red}0)\) again? ;)

OpenStudy (angelab97):

8

terenzreignz (terenzreignz):

That's right :) So we replace it with 8. \[\Large \frac{12\color{red}x+8}{8 }= 2\color{red}x\] Could you solve for x here? :)

OpenStudy (angelab97):

12x +8 = 16x 8=4x x=2

terenzreignz (terenzreignz):

Good job :)

OpenStudy (angelab97):

Thank you so much ! :))

terenzreignz (terenzreignz):

No problem :)